The loading dose of phenytoin in children is 20 mg/kg of body weight to be infused at a rate of 0.5 mg/kg/minute. Over how many minutes should the dose be administered to a 32 lb child?
Answer #1

Dose of phenytoin - 20mg/kg = 20× 14.4kg = 288mg

Flow rate - 0.5mg/kg/min = 0.5× 14.4kg= 7.2mg

Total time to infuse in min -?

Weight of the child - 32lbs = 32×0.45= 14.4kg

Flow rate = total dose /total time

= 288mg/7.2 mg /min

= 40min